# LENS: Learning to Segment Anything with Unified Reinforced Reasoning

### v1

@@ Introduction
Table 1: Comparison on the Referring Expression Segmentation (RES) task using cIoU as the evaluation metric. Our method achieves state-of-the-art performance across all RefCOCO, RefCOCO+, and RefCOCOg benchmarks.Bold indicates the best (SoTA) performance, and underlined denotes the second-best.
Method RefCOCO RefCOCO+ RefCOCOg AVG
val testA testB val testA testB val test
without active chain-of-thought (CoT) reasoning
LAVT (Yang et al. 2022)(CVPR 22) 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1 65.8
ReLA (Liu, Ding, and Jiang 2023)(CVPR 23) 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0 68.3
X-Decoder (Zou et al. 2023a)(CVPR 23) - - - - - - 64.6 - -
SEEM (Zou et al. 2023b)(NeurIPS 23) - - - - - - 65.7 - -
LISA (Lai et al. 2024) (CVPR 24) 74.1 76.5 71.1 62.4 67.4 56.5 66.4 68.5 67.9
PixelLM (Ren et al. 2024)(CVPR 24) 76.9 78.5 74.4 69.2 72.1 64.5 70.7 72.4 72.3
PerceptionGPT-7B (Pi et al. 2024)(CVPR 24) 75.1 78.6 71.7 68.5 73.9 61.3 70.3 71.7 71.4
OMG-LLaVA (Zhang et al. 2024a)(NeurIPS 24) 78.0 80.3 74.1 69.1 73.1 63.0 72.9 72.9 72.9
VISA (Yan et al. 2024)(ECCV 24) 72.4 75.5 68.1 59.8 64.8 53.1 65.5 66.4 65.7
GLaMM (Rasheed et al. 2024)(CVPR 24) 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9 75.6
with active chain-of-thought (CoT) reasoning
Seg-Zero-3B (Liu et al. 2025) (arXiv 25) - 79.3 - - 73.7 - 71.5 - -
Seg-Zero-7B (Liu et al. 2025) (arXiv 25) - 80.3 - - 76.2 - 72.6 - -
LENS-2B (Ours) 80.7 82.7 77.1 73.8 78.0 67.3 75.7 76.4 76.5
LENS-3B (Ours) 84.2 85.3 81.0 79.4 82.8 74.3 81.2 81.0 81.2
GRPO with Unified Rewards.
Unlike standard VLM-R1 methods (Shen et al. 2025; Yu et al. 2025) that relies on purely understanding reward signal, i.e., format reward and box IoU reward, we introduce a comprehensive multi-faceted reward system specifically designed for both understanding and segmentation. Our reward function encompasses three complementary components:
  • Box IoU Reward ($R_{\text{box}}$): Measures localization accuracy through IoU between predicted and ground truth bounding boxes. Same as the IoU metric, the range is $[0,1]$.
The unified reward function is formulated as:
where $\{\lambda_{i}\}_{i=1}^{3}$ are hyperparameters that balance different reward components.
Training Objective.
$\mathcal{J}_{\text{LENS}}(\theta)$ $=\mathcal{J}(\theta;R_{\text{unified}},\mathbb{D}_{\mathrm{KL}})$
$\quad+\alpha\mathcal{L}_{\text{seg}}(P_{\text{seg}},M_{\text{gt}})$ (12)
Experiment
Implementation Details
We employ Qwen-VL series models(Bai et al. 2025; Wang et al. 2024) as the reasoning model and SAM2-Large (Ravi et al. 2024) as the segmentation model. Specifically, LENS-2B is utilizes Qwen2-VL and LENS-3B uses Qwen2.5-VL. Training is conducted on 16 NVIDIA L40S GPUs, with our pipeline built upon the DeepSpeed engine (Rasley et al. 2020). Detailed experimental settings are summarized in the Appendix.
Benchmarks.
We evaluate the proposed framework on multiple benchmarks: Referring Expression Segmentation (RES) benchmarks (RefCOCO series) (Nagaraja, Morariu, and Davis 2016; Yu et al. 2016), GroundingSuite benchmarks (Hu et al. 2025), and Reasoning Segmentation benchmarks (Lai et al. 2024). The quantitative comparisons are presented in Table 1 and Table 2, respectively.
Evaluation Metrics.
Following previous works on referring segmentation (Kazemzadeh et al. 2014; Yu et al. 2016), we adopt two commonly used evaluation metrics: generalized IoU (gIoU) and cumulative IoU (cIoU). Specifically, gIoU is computed as the average of per-image Intersection-over-Union (IoU) scores, while cIoU is defined as the IoU of the cumulative predicted and ground-truth masks across the entire dataset.
Main Results
Table 2: Performance comparison on the ReasonSeg and GroundingSuite-Eval benchmarks, evaluated using gIoU and cIoU. Bold and underlined entries indicate the best and second-best results, respectively. $\ddagger$ marks models trained on both train and test splits of ReasonSeg. Our approach shows notable superiority in cIoU across both datasets.
Method ReasonSeg-Val ReasonSeg-Test GS-Eval
gIoU cIoU gIoU cIoU gIoU cIoU
PSALM (Zhang et al. 2024c)(ECCV 24) - - - - 39.2 34.4
[…]
LISA-7B(ft) (Lai et al. 2024) (CVPR 24) 52.9 54.0 55.6 56.9 60.9 68.6
Seg-Zero-3B (Liu et al. 2025) (arXiv 25) 58.2 53.1 56.1 48.6 67.3 68.9
LENS-3B (Ours) 62.1 64.9 57.2 58.0 67.0 78.3
Note that PSALM (Zhang et al. 2024c), HyperSeg (Wei et al. 2024a), and InstructSeg (Wei et al. 2024b) utilize pre-trained Mask2Former (Cheng et al. 2022) models for mask generation, whose training data, i.e., COCO 2017 (Lin et al. 2014), has overlap with the RefCOCO benchmarks. Therefore, we compare with them in Table 2, evaluating their performance on the Reasoning Segmentation and GroundingSuite-Eval benchmarks.
Query Num. 1 16 32 64 128
cIoU (%) 85.9 87.4 87.8 87.9 87.2
Connector MLP ViT
cIoU (%) 72.8 87.8
Table 5: Ablation study of components on RefCOCO testA.
Context Module Align. Stage RL Stage cIoU
$R_{\text{box}}$ $R_{\text{format}}$ $R_{\text{seg}}$
$\times$ $\times$ $\times$ $\times$ $\times$ 79.3
$\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ 72.3
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 81.9
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 82.7
Referring Expression Segmentation.
As shown in Table 1, the proposed LENS achieves state-of-the-art performance across all splits of the RefCOCO, RefCOCO+, and RefCOCOg benchmarks. Compared to prior models such as GlaMM (Rasheed et al. 2024), our method delivers significant improvements, especially on RefCOCO+ and RefCOCOg, which are known to require a more precise understanding of compositional and spatial language. Notably, LENS-3B surpasses the best previous method by 6.1% on RefCOCOg-test, highlighting the model’s strength in handling complex referring expressions.
ReasonSeg and GroundingSuite-Eval Benchmarks.
Table 2 presents results on the ReasonSeg and GroundingSuite-Eval benchmarks. The proposed LENS achieves competitive performance in Generalized IoU (gIoU), and clearly leads in cumulative IoU (cIoU), which reflects more holistic segmentation quality across objects and scenes. For example, the proposed LENS achieves a 78.3% cIoU on GroundingSuite-Eval, outperforming Seg-Zero-3B by 9.4%.
Ablation Studies
Number of Queries.
Connector Design.
We further investigate the role of the proposed connector that bridges the MLLM and the segmentation head. As shown in Table 4, replacing a simple MLP with a Vision Transformer (ViT) module significantly boosts performance. Overall, these studies underscore the importance of architectural choices in enabling effective reasoning and accurate segmentation within the proposed unified framework.
Framework Analysis.
To investigate the contribution of each component in our framework, we conduct an ablation study on the RefCOCO testA split, as shown in Table 5. We start with a Lisa-liked baseline (Lai et al. 2024) model without the context module, alignment stage, or reinforcement learning (RL), which achieves a cIoU of 79.3%. When directly adding the context module and applying RL with the box reward, the model can not stably following our instructions and output reasoning contexts. After enabling format reward, the model captures basic localization cues, but the lack of explicit alignment between textual and visual features leads to sub-optimal performance. To address this issue, we enable the alignment stage, where the context module is pretrained to align MLLM and SAM. This significantly improves the segmentation accuracy, boosting cIoU to 81.9%, and demonstrates the critical importance of semantic alignment in the proposed unified framework. Finally, introducing the segmentation-level reward ($R_{\text{seg}}$) further refines the mask prediction quality, pushing the performance to 82.7% cIoU.
Visualizations
We provide qualitative results to illustrate the effectiveness of our framework on both general and reasoning-intensive segmentation tasks. Fig. 3 shows examples from the RefCOCOg benchmark. LENS accurately segments target objects based on referring expressions, demonstrating strong performance in scenes with partially obscure objects and multiple similar instances. Notably, the proposed unified framework and end-to-end optimization make LENS robust to correct the potential error in the box prior. Fig. 4 presents results on the reasoning segmentation benchmark. The model exhibits robust test-time reasoning capabilities, effectively handling tasks that require spatial understanding, multi-step inference, and comparative reasoning. These visualizations highlight the model’s ability to integrate high-level textual reasoning with fine-grained visual perception across diverse scenes.
Conclusion
Appendix A Appendices
AAAI 2026 Supplementary Material
Anonymous Submission
Anonymous Submission 1School of EIC, Huazhong University of Science & Technology, 2vivo AI Lab
Code & Models: hustvl/LENS
Appendix B Implementation Details
@@
For the pretraining alignment stage, we follow the settings in SAM2 (Ravi et al. 2024) and Qwen2.5-VL (Bai et al. 2025). Moreover, we list the important experimental settings as shown in Table A1. Note that the training set in this stage is from the RefCOCO series (Nagaraja, Morariu, and Davis 2016; Yu et al. 2016), which includes RefCOCO, RefCOCO+, and RefCOCOg.
Table A1: Pretraining alignment stage experimental settings. We use AdamW (Loshchilov and Hutter 2017) as the optimizer.
@@
For the referring expression segmentation, we follow the settings in VLM-R1 (Shen et al. 2025) and list the important experimental settings as shown in Table A2.
Referring Expression Segmentation
Table A2: Reinforcement learning stage experimental settings. We use AdamW (Loshchilov and Hutter 2017) as the optimizer.
@@
For the referring expression segmentation, our training set is from the RefCOCO series (Nagaraja, Morariu, and Davis 2016; Yu et al. 2016), including RefCOCO, RefCOCO+, and RefCOCOg.
Reasoning Segmentation
For the reasoning segmentation, we adopt the ReasonSeg (Lai et al. 2024) for training.
Appendix C More Visualization Examples
@@
As shown in Fig. A1, we showcase examples where the model successfully interprets spatial cues in the referring expressions, such as ”third bottle from the left” or ”the orange in the 5 o’clock position”. These cases highlight the model’s ability to integrate language-based spatial relations with visual layout.
Figure A1: Example of spatial understanding. Our method correctly segments the target object based on spatial cues in the referring expression.
@@
As shown in Fig. A2, we present examples in which multiple similar objects are present in the scene, and the referring expression must disambiguate between them. Our method is able to correctly segment the intended target by leveraging distinguishing attributes in the language input.
Figure A2: Example of disambiguation among multiple similar objects. Our method recognizes distinguishing attributes in the referring expression and accurately identifies the intended target.
@@
Figure A3: Example of handling complex referring expressions. Our method can process multiple attributes or relationships in the referring.
@@
Figure A4: Example of segmentation under misguided context. When the initial context box is incorrect or ambiguous, our method can leverage multi-query context embeddings to generate an accurate segmentation mask, demonstrating robustness to context errors.
@@
Figure A5: Example of coarse annotations. Our method predicts a mask that captures fine-grained object boundaries and surpasses the coarse or imprecise ground-truth mask annotations.
@@
As shown in Fig. A6, we further visualize samples that require reasoning beyond simple visual matching, such as temporal states (e.g., ”the person who is eating”) or implicit object relationships. The results demonstrate the model’s capacity for complex reasoning segmentation.
Figure A6: Example of reasoning segmentation. Our method successfully infers and segments the target based on a complex instruction.

### v2

@@ Appendix A Appendices
More Implementation Details
@@
For the pretraining alignment stage, we follow the settings in SAM2 (ravi2024sam2segmentimages) and Qwen2.5-VL (bai2025qwen2p5vl). Moreover, we list the important experimental settings as shown in Table A1. Note that the training set in this stage is from the RefCOCO series (nagaraja2016modeling; yu2016modeling), which includes RefCOCO, RefCOCO+, and RefCOCOg.
@@
Table A1: Pretraining alignment stage experimental settings. We use AdamW (loshchilov2017decoupled) optimizer.
@@
For the referring expression segmentation, we follow the experimental settings in VLM-R1 (shen2025vlmr1) as shown in Table A2.
@@
Table A2: Reinforcement learning stage experimental settings. Optimizer is AdamW (loshchilov2017decoupled).
For the referring expression segmentation, our training set is from the RefCOCO series, including RefCOCO, RefCOCO+, and RefCOCOg. For the reasoning segmentation, we adopt the ReasonSeg (lai2024lisa) for training.
More Ablations
SAM Version Tune SAM Context Module Align. Stage RL Stage cIoU 1 $\times$ $\times$ $\times$ $\times$ 78.5 1 $\checkmark$ $\times$ $\times$ $\times$ 79.1 2 $\checkmark$ $\times$ $\times$ $\times$ 79.3 2 $\checkmark$ $\checkmark$ $\times$ $\times$ 80.4 2 $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 80.9 2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 82.7
Table A3: Extended ablation study on SAM variants and training strategy on RefCOCO testA.
MLLM Context Query Connector SAM2 Params (M) 3755.6 0.13 206.3 224.4 Latency (s) 1.5 0.07 0.002 0.025
Table A4: Parameter and latency of model components.
We conduct an additional ablation on RefCOCO testA to examine the impact of SAM variants and training choices (Table A3). Replacing SAM2 by SAM1 or freezing the mask decoder reduces performance, which highlights the benefit of a stronger SAM backbone and fine-tuning the mask decoder. Adding the context module improves cIoU to 80.4% by capturing sentence-level and cross-modal cues. Introducing the Alignment Stage, which explicitly aligns textual and visual features, further increases performance to 80.9%, demonstrating the importance of semantic alignment between the MLLM and SAM. Finally, incorporating the segmentation-level RL stage refines mask quality and yields the largest gain, reaching 82.7% cIoU. Overall, these results indicate that a stronger off-the-shelf segmenter, context modeling, explicit alignment, and RL each provide complementary benefits and collectively lead to substantial improvements.
Rewards and Loss During Training
The GRPO training process demonstrates stable and effective policy learning, as evidenced by the concurrent trends in reward and loss (Fig. A1). We define the IoU reward as the squared segment IoU and the squared box IoU, i.e., reward is $\text{IoU}_{\text{seg}}^{2}$ or $\text{IoU}_{\text{box}}^{2}$. We observe a steady increase in both format reward, segment IoU, and box IoU rewards, which correlates with the generation of better-formatted responses, higher-quality masks and more accurate grounding predictions. The reward improvements corresponds with a steady decrease in training loss, confirming the convergence of the optimization process.
More Module Details
Our multimodal large language model (MLLM) is Qwen2.5-VL-3B, whose embedding dimension is 2048. The context query module consists of 64 learnable embeddings, which serve as task-specific query slots for extracting and structuring multimodal reasoning signals from the MLLM. The connector module is implemented as a 4-layer Qwen2.5-style Transformer with a hidden dimension of 2048, ensuring dimensional alignment with the MLLM outputs. It maps textual and segmentation prompts into a unified embedding space, enabling seamless interaction between the MLLM’s high-level reasoning and the SAM2 segmentation head.
As summarized in Tab. A4, the context query and connector together contribute only 4.9% of the total parameters and 4.5% of the overall inference latency. Despite being lightweight, these modules provide an effective and efficient bridge between multimodal understanding and mask generation, significantly enhancing grounding and segmentation performance with minimal computational overhead.
More Visualization Examples
@@
As shown in Fig. A2, we showcase examples where the model successfully interprets spatial cues in the referring expressions, such as “third bottle from the left” or “the orange in the 5 o’clock position”. These cases highlight the model’s ability to integrate language-based spatial relations with visual layout.
Figure A1: Over Unified GRPO training, the format reward, segment reward, box reward, and the loss curves indicate steady reward gains and consistent loss decreases.
Figure A2: Example of spatial understanding. LENS correctly segments the target object based on spatial cues in the referring expression.
@@
As shown in Fig A3, we present examples in which multiple similar objects are present in the scene, and the referring expression must disambiguate between them. Our method is able to correctly segment the intended target by leveraging distinguishing attributes in the language input.
Figure A3: Example of disambiguation among multiple similar objects. LENS recognizes distinguishing attributes in the referring expression and accurately identifies the intended target.
@@
Figure A4: Example of handling complex referring expressions. LENS can process multiple attributes or relationships in the referring.
@@
Figure A5: Example of segmentation under misguided context. When the initial context box is incorrect or ambiguous, LENS can leverage multi-query context embeddings to generate an accurate segmentation mask, demonstrating robustness to context errors.
@@
Figure A6: Example of coarse annotations. LENS predicts a mask that captures fine-grained object boundaries and surpasses the coarse or imprecise ground-truth mask annotations.
@@
As shown in Fig. A7, we further visualize samples that require reasoning beyond simple visual matching, such as temporal states (e.g., “the person who is eating”) or implicit object relationships. The results demonstrate the model’s capacity for complex reasoning segmentation.
Figure A7: Example of reasoning segmentation. LENS successfully infers and segments the target based on a complex instruction.
More Discussion about Failure Cases
Our method may fail in cases where the instruction requires fine-grained reasoning, refers to ambiguous attributes, or when the target object is very small. Fig. A8 summarizes several representative failure modes. (1) Granularity confusion: when the instruction refers to “a plate with fruit on it,” our model segments both the plate and the fruit, whereas the ground truth only annotates the plate. This arises from the lack of part-level supervision. (2) Wrong-instance selection: for the instruction “the clock showing 12:21,” the model selects a different clock instance with mismatched attributes. (3) Annotation error: in “a knife cutting a cake,” the ground truth incorrectly labels the person holding the knife instead of the knife itself. (4) Small object difficulty: for “the chair the bear on the right is sitting in,” the target chair is very small in the image, which makes precise segmentation challenging and can lead to over-segmentation.
Figure A8: Failure cases: (1) granularity confusion, (2) wrong-instance selection, (3) annotation error, and (4) small object difficulty.